# Prof. Robot: Differentiable Robot Rendering Without Static and Self-Collisions

### v1

@@ Abstract
Equal contribution.* Corresponding authors.
@@ 3 Method
Figure 2: Our Pipeline. (Left) We predict whether a collision has occurred based on pose input, using Eikonal loss ($\mathcal{L}_{\textrm{ek}}$) and binary cross-entropy loss ($\mathcal{L}_{\textrm{BCE}}$) to ensure this. It can then be seamlessly integrated with Dr. Robot [32] to create a fully differentiable pipeline for optimization.
@@ 3.1 Collision Detection by Pose Classification
$\mathcal{V}=\left\{\bm{\theta}\in\textrm{SO}(2)^{K}\mid f(\bm{\theta})\leq 0.5\right\}.$
@@
$\mathcal{L}_{\textrm{BCE}}=\textrm{BCE}(f(\bm{\theta}),c^{\textrm{gt}}),$
@@ 3.2 Eikonal Regularizer for Consistent Gradients
$\left\langle\frac{\partial g}{\partial\bm{\theta}_{t}},\frac{\partial g}{\partial\bm{\theta}_{t+1}}\right\rangle\leq 0,$
@@
$\left\langle\frac{\partial g}{\partial\bm{\theta}_{t}},\frac{\partial g}{\partial\bm{\theta}_{t+1}}\right\rangle\approx\left\langle\frac{\partial g}{\partial\bm{\theta}_{t}},\frac{\partial g}{\partial\bm{\theta}_{t}}\right\rangle=\left\|\frac{\partial g}{\partial\bm{\theta}_{t}}\right\|^{2}\geq 0.$
@@
$\mathcal{L}_{\textrm{ek}}=\left(\left\|\frac{\partial g}{\partial\bm{\theta}}\right\|-a\right)^{2}$
@@
$f(\bm{\theta})=(\sigma\circ\omega\circ g)(\bm{\theta}),$
@@ 3.4 Training Data and Objectives
$\mathcal{L}=\mathcal{L}_{\textrm{BCE}}+\alpha\mathcal{L}_{\textrm{ek}}.$
@@ 4.3 Collision Resolution by Pose Optimization
4.4 Integration with differentiable robot rendering
Differentiable robot rendering enables us to obtain a differentiable pose from the rendered view. From this pose, our method can infer prior knowledge about potential collisions with the external environment or self-intersections. This prior knowledge can be seamlessly integrated into the differentiable rendering process.
@@ 4.4.1 Inverse action from collision image
$\mathcal{L}_{\textrm{RGB}}=\frac{1}{N}\sum_{\textrm{seg}\left[i,j\right]=1}\|\bm{r}-\bm{r}_{\textrm{gt}}\|$
$\mathcal{L}_{\textrm{D}}=\frac{1}{N}\sum_{\textrm{seg}\left[i,j\right]=1}\|\bm{d}-\bm{d}_{\textrm{gt}}\|$
@@
$\mathcal{L}=\mathcal{L}_{\textrm{RGB}}+\mathcal{L}_{\textrm{D}}$
@@
$\mathcal{L}_{\textrm{SDF}}=\textrm{ReLU}\left(g\left(\bm{\theta}\right)-l_{\textrm{thres}}\right)$
@@
$\mathcal{L}=\mathcal{L}_{\textrm{RGB}}+\mathcal{L}_{\textrm{D}}+\beta\cdot\mathcal{L}_{\textrm{SDF}}$
@@ 4.4.2 Inverse Trajectory from Optimization
Figure 5: Control Experiment. The red frame represents the manipulation from Dr. Robot, whereas the green frame highlights ours. A heatmap has been utilized to visually demonstrate the collision between the robot and the desk. Our method allows for the stable pick of the gray block from the desk.
@@
As mentioned in Section 4.4.1, we randomly sample two collision-free poses and obtain their respective rendering results. Dr. Robot uses $\bm{\theta}_{\textrm{start}}$ as the initial configuration and utilizes the image, mask, depth, and pose information of $\bm{\theta}_{\textrm{end}}$ from the simulator as supervisory signals. In addition, we incorporate our SDF model and add $\mathcal{L}_{\textrm{SDF}}$ to the loss function.
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\textrm{SDF}}\left(\hat{\bm{\theta}_{i}}\right)+\frac{1}{N-1}\sum_{i=1}^{N-1}\|\hat{\bm{\theta}}_{i+1}-\hat{\bm{\theta}_{i}}\|$
@@ 4.4.3 Inverse Action from VLM
4.4.4 Control Robot from Video Input
The outcomes of our experiments are delineated in Table 10 and illustrated in Figure 5. The data reveals that initially, the actions derived from Dr. Robot resulted in numerous collisions. However, upon implementing our optimization technique, it became evident that collisions could be significantly mitigated. This underscores the success of our method in enhancing the robot’s operational safety.
Table 10: The metrics of collision during robot manipulation.
Collision Point $\downarrow$ Collision Rate $\downarrow$
Dr. Robot 0.62 15.71%
Ours 0.00 0.00%
@@ 4.5 Sim-to-Real Deployment
4.6 Comparison with Differentiable Simulator
4.7 Train with Alternative Regularizers
@@ 5 Limitations and Future Work
Our method face to the problem of cross-environment generalization. To tackle this, a new mechanism that enhances Prof. Robot’s scene perception through dynamic object modeling could be proposed. The dynamic environment could be represented by supplementary joints $\delta$, which constrain the robotic system’s joint space. By training with a collision-aware discriminator $p(\cdot|\delta)$, our joint learning framework could practical effectiveness despite environmental perturbations from ideal conditions. In practical scenarios, this is achieved by estimating the poses of dynamic objects.
@@ 8.1 Trajectory of Interpolation
$\hat{\bm{\theta}_{1}},\ldots,\hat{\bm{\theta}_{N}}=\mathcal{I}\left(\bm{\theta}_{\textrm{start}},\bm{\theta}_{\textrm{end}}\right)$
@@
$\mathcal{L}=\gamma_{1}\mathcal{L}_{\textrm{SDF}}+\gamma_{2}\mathcal{L}_{\textrm{TV}}$
@@
$\mathcal{L}_{\textrm{TV}}=\frac{1}{N-1}\sum_{i=1}^{N-1}\|\hat{\bm{\theta}}_{i+1}-\hat{\bm{\theta}}_{i}\|$
@@
$\mathcal{L}=\gamma_{1}\mathcal{L}_{\textrm{SDF}}+\gamma_{2}\mathcal{L}_{\textrm{TV}}+\gamma_{3}\mathcal{L}_{\textrm{SDF}}$
@@ 8.3 Trajectory of Tangent
@articlenewbury2024review, title=A review of differentiable simulators, author=Newbury, Rhys and Collins, Jack and He, Kerry and Pan, Jiahe and Posner, Ingmar and Howard, David and Cosgun, Akansel, journal=IEEE Access, year=2024, publisher=IEEE

### v2

@@ Abstract
Equal contribution.* Corresponding authors.  Project: https://qrcat.github.io/prof-robot/
@@ 3 Method
Figure 2: Our Pipeline. We predict whether a collision has occurred based on pose input, using Eikonal loss ($\mathcal{L}_{\textrm{ek}}$) and binary cross-entropy loss ($\mathcal{L}_{\textrm{BCE}}$) to ensure this. It can then be seamlessly integrated with Dr. Robot [32] to create a fully differentiable pipeline for optimization.
@@ 3.1 Collision Detection by Pose Classification
$\mathcal{V}=\left\{\bm{\theta}\in\textrm{SO}(2)^{K}\mid f(\bm{\theta})\leq 0.5\right\}.$ (1)
@@
$\mathcal{L}_{\textrm{BCE}}=\textrm{BCE}(f(\bm{\theta}),c^{\textrm{gt}}),$ (2)
@@ 3.2 Eikonal Regularizer for Consistent Gradients
$\left\langle\frac{\partial g}{\partial\bm{\theta}_{t}},\frac{\partial g}{\partial\bm{\theta}_{t+1}}\right\rangle\leq 0,$ (3)
@@
$\left\langle\frac{\partial g}{\partial\bm{\theta}_{t}},\frac{\partial g}{\partial\bm{\theta}_{t+1}}\right\rangle\approx\left\langle\frac{\partial g}{\partial\bm{\theta}_{t}},\frac{\partial g}{\partial\bm{\theta}_{t}}\right\rangle=\left\|\frac{\partial g}{\partial\bm{\theta}_{t}}\right\|^{2}\geq 0.$ (4)
@@
$\mathcal{L}_{\textrm{ek}}=\left(\left\|\frac{\partial g}{\partial\bm{\theta}}\right\|-a\right)^{2}$ (5)
@@
$f(\bm{\theta})=(\sigma\circ\omega\circ g)(\bm{\theta}),$ (6)
@@ 3.4 Training Data and Objectives
$\mathcal{L}=\mathcal{L}_{\textrm{BCE}}+\alpha\mathcal{L}_{\textrm{ek}}.$ (8)
@@ 4 Experiments and Results
We conduct ablation studies on Prof. Robot (section 4.2), experiments on collision resolution (section 4.3), integration with Dr. Robot (section 4.4), and sim-to-real deployment (section 4.5). Comparison with the differentiable simulator, training with alternative regularizers, and manipulation with Dr. Robot are detailed in the supplementary materials.
@@ 4.3 Collision Resolution by Pose Optimization
4.4 Integration with Dr. Robot
Differentiable robot rendering (Dr. Robot) enables us to obtain a differentiable pose from the rendered view. From this pose, our method can infer prior knowledge about potential collisions with the external environment or self-intersections. This prior knowledge can be seamlessly integrated into the differentiable rendering process.
@@ 4.4.1 Inverse Action from Collision Image
$\mathcal{L}_{\textrm{RGB}}=\frac{1}{N}\sum_{\textrm{seg}\left[i,j\right]=1}\|\bm{r}-\bm{r}_{\textrm{gt}}\|$ (9)
$\mathcal{L}_{\textrm{D}}=\frac{1}{N}\sum_{\textrm{seg}\left[i,j\right]=1}\|\bm{d}-\bm{d}_{\textrm{gt}}\|$ (10)
@@
$\mathcal{L}=\mathcal{L}_{\textrm{RGB}}+\mathcal{L}_{\textrm{D}}$ (11)
@@
$\mathcal{L}_{\textrm{SDF}}=\textrm{ReLU}\left(g\left(\bm{\theta}\right)-l_{\textrm{thres}}\right)$ (12)
@@
$\mathcal{L}=\mathcal{L}_{\textrm{RGB}}+\mathcal{L}_{\textrm{D}}+\beta\cdot\mathcal{L}_{\textrm{SDF}}$ (13)
@@ 4.4.2 Inverse Trajectory from Optimization
As mentioned in Section 4.4.1, we randomly sample two collision-free poses and obtain their respective rendering results. Dr. Robot uses $\bm{\theta}_{\textrm{start}}$ as the initial configuration and utilizes the image, mask, depth, and pose information of $\bm{\theta}_{\textrm{end}}$ from the simulator as supervisory signals. In addition, we incorporate our SDF model and add $\mathcal{L}_{\textrm{SDF}}$ (Equation 12) to the loss function.
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\textrm{SDF}}\left(\hat{\bm{\theta}_{i}}\right)+\frac{1}{N-1}\sum_{i=1}^{N-1}\|\hat{\bm{\theta}}_{i+1}-\hat{\bm{\theta}_{i}}\|$ (14)
@@ 5 Limitations and Future Work
Our method tackles the problem of cross-environment generalization. To address this, a new mechanism could be proposed that enhances Prof. Robot’s scene perception by modeling dynamic objects. The dynamic environment could be represented by supplementary joints $\delta$, which constrain the robotic system’s joint space. By training with a collision-aware discriminator $p(\cdot|\delta)$, our joint learning framework could practical effectiveness despite environmental perturbations from ideal conditions. In practical scenarios, this is achieved by estimating the poses of dynamic objects.
@@ 8.1 Trajectory of Interpolation
$\hat{\bm{\theta}_{1}},\ldots,\hat{\bm{\theta}_{N}}=\mathcal{I}\left(\bm{\theta}_{\textrm{start}},\bm{\theta}_{\textrm{end}}\right)$ (15)
@@
$\mathcal{L}=\gamma_{1}\mathcal{L}_{\textrm{SDF}}+\gamma_{2}\mathcal{L}_{\textrm{TV}}$ (16)
@@
$\mathcal{L}_{\textrm{TV}}=\frac{1}{N-1}\sum_{i=1}^{N-1}\|\hat{\bm{\theta}}_{i+1}-\hat{\bm{\theta}}_{i}\|$ (17)
@@
$\mathcal{L}=\gamma_{1}\mathcal{L}_{\textrm{SDF}}+\gamma_{2}\mathcal{L}_{\textrm{TV}}+\gamma_{3}\mathcal{L}_{\textrm{SDF}}$ (18)
@@ 8.3 Trajectory of Tangent
9 Comparison with Differentiable Simulator
10 Train with Alternative Regularizers